OpenStudy (anonymous):

Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m2 by increasing the width and length by the same amount. What will be the width (shorter dimension) of the new garden? A. 13 m wide B. 14 m wide C. 15 m wide D. 16 m wide

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

can it be to answers?

OpenStudy (paki):

solve for "x"....

OpenStudy (anonymous):

it is either A or D

OpenStudy (paki):

D is the right answer...

OpenStudy (anonymous):

AHA! its ........ A

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its A because you need to add the same number

OpenStudy (paki):

coolnidal just make an equation... and then solve it... you will get 16 sir :-)

OpenStudy (anonymous):

add 3 to original width and to the original length equal the answer a

OpenStudy (anonymous):

no i am sure its A

OpenStudy (anonymous):

both of you are wrong its c but thank you for trying

OpenStudy (anonymous):

the answer choices are for width only

OpenStudy (anonymous):

wt!

OpenStudy (anonymous):

thank you lol

OpenStudy (paki):

at then end you will get... 208/13-10=x x=208-13/13 x=78/13 x=6 :-) pleasure @sierraleone17